Internal Rate of Return (IRR) can be understood as the discount rate that should be applied...

  1. Internal Rate of Return (IRR) can be understood as the discount rate that should be applied to a project such that Net Present Value (NPV) = $0. If the discount rate applied in a certain 5 year project is 10%, the resultant NPV is $50K, and the cost is $100K, what is the IRR? (Assume that all costs for the project are incurred at the start of the project, and the payout for this project occurs in one payment at the end of Year 5).
Answer #1

When interest rate of 10% is considered then NPV= Pv of Benefits-PV of Costs

Now NPV=$50k & PV of Cost=$100 ( because we have incurred cost at the beginning)

50=B/(1.1)^5-100

B/(1.1)^5=150

B=150(1.1^5)=241.58

IRR is r when

241.58/(1+r)^5=100

2.4158=(1+r)^5

2.4158^0.2-1=r

r=19.29% is the internal rate of return for this project